13. A firm has hired a management consulting firm to analyze demand for its product. The results relating to this demand function are presented below: Qx = 1572 + 2A + 10Y + 3PY – 4Px Where, Qx = quantity demand for x good A = advertisement expenditure in thousands P = price of x good in rupees Py = price of another product in rupees Y = income in thousands a. Interpret the constant parameter and coefficients corresponding to Y and Px in the given demand equation. b. Find the quantity demanded if P =Rs 400, Y = Rs 15000, A = Rs 20000 and Py = Rs 500 c. If price is reduced to Rs 300, what is the change in quantity demand? d. Is the demand elastic or inelastic? What would be the effect of a rise in the price on the total income? e. Are two goods substitutes or complements? Give reason. 


1
Expert's answer
2022-02-28T09:38:54-0500

a) The constant value is constant elasticity of demand and it is buffer zone units of production.

-4 is elasticity and 10 measures the effect of unit change in demand on income.


b) Qx=1572 + 2*20000 + 10*15000 + 3*500 – 4*400

Qx=1572+40000+150000+1500-1600

Qx=56,472 units


c) Qx=1572+40000+150000+1500-1200

Qx=56,872 units; quantity demanded will increase by 400 units


d) Elastic because price changes affect demand. Rise in price will have no effect on total income.


e) The two goods are compliments because, they prohibit the same features and they have a constant that is positive.